the area of a rectangle is 165ft^2. the length is 4 ft greater than the width.
find the dimension of the rectangle.
w=
l=

Answer by uday1435(57)

l = w+4
Area = lw = (w+4)w = w^2 + 4w = 165
w^2 + 4w – 165 = 0 factorizing we get
(w+15)(w – 11) = 0
So the width = 11 ft
And length = 11+4 = 15 ft


If you still require further clarifications please write to udayakumar.t.r@gmail.com